# Not Everyone Wins with LLMs: Behavioral Patterns and Pedagogical Implications in AI-assisted Data Analysis

### v1

@@ Abstract.
LLMs promise to democratize technical work in complex domains like programmatic data analysis, but not everyone benefits equally. We study how students with varied expertise use LLMs to complete Python-based data analysis in computational notebooks in a non-major course. Drawing on homework logs, recordings, and surveys from 36 students, we ask: Which expertise matters most, and how does it shape AI use? Our mixed-methods analysis shows that technical expertise – not AI familiarity or communication skills – remains a significant predictor of success. Students also vary widely in how they leverage LLMs, struggling at stages of forming intent, expressing inputs, interpreting outputs, and assessing results. We identify success and failure behaviors, such as providing context or decomposing prompts, that distinguish effective use. These findings inform AI literacy interventions, highlighting that lightweight demonstrations improve surface fluency but are insufficient; deeper training and scaffolds are needed to cultivate resilient AI use skills.
large language model, data science assistant, data analysis, AI coding assistant, human-AI interaction
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXbooktitle: 2025 Preprintprice: 15.00isbn: 978-X-XXXX-XXXX-X/YY/MMccs: Human-centered computing Interactive systems and toolsccs: Applied computing Computer-assisted instructionccs: Computing methodologies Natural language generation
@@ 1. Introduction
Large language models (LLMs) hold the promise of democratizing complex technical tasks such as data analysis by assisting users in programming, analysis, and modeling (Ma et al., 2025; Pickering et al., 2025). Despite the grand vision, we continue to see mixed results in whether and how people benefit from AI (Paradis et al., 2025). There have been conflicting reports on whether AI agents help developers improve programming efficiency (Becker et al., 2025; Song et al., 2024); only some novices seem to perceive reduced barriers (Pickering et al., 2025; Kazemitabaar et al., 2023). In fact, performance often hinges on users’ prior expertise, ranging from domain knowledge to AI literacy (Chen et al., 2024; Nguyen et al., 2024a; Zi et al., 2025). For instance, some work argues that less experienced developers benefit more from AI assistants (Liang et al., 2024; Peng et al., 2023), and those with basic AI literacy may avoid common pitfalls in prompting or interpreting model outputs (Puppart and Aru, 2025; Annapureddy et al., 2025). However, these findings remain fragmented across different populations and tasks, making it challenging to weigh which types of expertise matter most (e.g., whether we should focus on teaching domain knowledge or AI usage knowledge). Moreover, most prior analyses measure only outcomes (e.g., task success) and are conducted in controlled, short-term lab settings (Kazemitabaar et al., 2024a; Pickering et al., 2025; Zheng, 2023). As a result, we have yet to demystify how users’ AI interaction behaviors get shaped by their (lack of) certain expertise in real-world workflows.
@@
In this work, we aim to examine how varied forms of expertise influence AI use and outcomes in a naturalistic classroom setting: a graduate-level course, Data Science for Product Management (DSPM). The class offered a unique opportunity to study LLM-supported workflows: it included a diverse cohort of 40 students from CS majors to business students transitioning into technical roles, allowing us to examine varying domain expertise, LLM literacy, and communication skills. Students worked independently on a series of data analysis assignments (e.g., Data Cleaning) in Google Colab, had full access to the built-in Gemini LLM assistant (Figure 1), and their activity was captured through logs, surveys, and screen recordings. Through this setup, we investigate three questions of increasing depth:
RQ1 (Section 5): Do LLMs close the performance gap between students with different expertise? We first conduct quantitative analysis on the relationship between students’ self-perceived expertise (in technical, LLM, and communication domains) and their grades across four assignments. Despite all students having equal access to Gemini, we find that self-reported technical expertise significantly predicted student grades, while LLM expertise and communication skills did not. This suggests that LLMs do not fully close the expertise gap — students with stronger programming and data science backgrounds outperform their peers when using AI. Many dimensions of expertise were highly correlated in our study, making it difficult to isolate their independent contributions. This indicates that fully isolating expertise may require more deliberate sampling and experimental control in future work.
@@
RQ2 (Section 6): How do differences in expertise translate to AI usage behaviors? We analyze students’ interaction logs using a novel behavior-grounded codebook, developed via thematic analysis and scaled up through LLM-based annotation. Our method segments over 7,000 raw log events into intent-driven episodes and steps, where each step is annotated with the presence (or absence) of effective AI behaviors, such as, decomposing tasks, providing contexts, and using AI for explanation or critique. Our analysis shows that technical experts use AI more strategically. They write clearer prompts, are more likely to provide necessary context, and use AI proactively to explore alternative solutions or improve plans (e.g., asking for better visualizations). In contrast, novices rely more on AI to resolve immediate challenges, such as debugging errors or explaining confusing outputs. These behavioral patterns offer a more nuanced explanation of why expertise continues to matter, even in AI-supported environments.
@@
RQ3 (Section 7): What and how can we train students on AI usage? We analyze missed opportunities for AI use by identifying log steps where AI could have been helpful but was not used. From this, we distilled a set of AI-use competencies spanning across four knowledge dimensions Conceptual, Procedural, Metacognitive, Dispositional. We show that some AI-use skills can be improved via live coding demonstration and extended time-on-task (e.g., regarding prompt quality), while others likely require guided practice and feedback (e.g., regarding evaluation), and are hard to cover via demonstration (e.g., metacognitive and dispositional skills). In sum, we make three main contributions:
  • Empirical evidence on expertise vs. AI use: We show that even in open-ended, LLM-supported workflows, technical expertise remains a significant predictor of success, and self-reported communication skill and AI familiarity alone do not suffice.
  • Complex log-based characterization of AI use: We introduce a behavior-oriented, LLM-based annotation schema that segments over 7,000 raw log events into semantically meaningful user stories. Compared to prior work that analyzes linear chat logs (Lee et al., 2022), our setup in Google Colab involves multiple LLM-powered functions with different affordances and model access, resulting in more nuanced observations about when, how, and why students turn to AI support (e.g., comparing behaviors between AI-as-chat versus AI-as-inline-tool). We open-source all of our scripts, including log data collection, cleaning, annotation, and analysis: https://github.com/mqo00/dspm
  • Pedagogical implications: We identify which AI skills can be improved via lightweight instruction vs. require deeper pedagogical investment. This informs future curricula and tool design for training effective AI collaborators.
@@ 2. Related Works
2.1. Novice versus Expert Strategies in Programming with AI
HCI and educational research highlight systematic differences between novices and experts when collaborating with AI for programming and modeling. In agent-based modeling with ChatGPT, experts benefit more than novices and rely on stronger skills in task decomposition and evaluation of AI outputs (Chen et al., 2024). In data-science code comprehension, experts inspect the data deeply, verify assumptions about transformations, and navigate outputs with clear goals—behaviors novices underuse (Lum et al., 2025). At the same time, studies document failure patterns such as “prompting rabbit holes,” where users iterate on prompts without convergence (Tie et al., 2024). However, most prior studies also focus on challenges brought by domain expertise differences, such as experts verifying and examining the data more deeply, and few focus on domain-agnostic LLM use patterns or skills regardless of expertise (Section 2.3). Together, these findings motivate our research questions: measure whether LLMs truly reduce expertise gaps, and analyze which behaviors distinguish effective from ineffective AI use in authentic notebook workflows.
@@ 2.2. Interfaces and Log Mining for AI-Assisted Data Analysis
AI-based computational notebook assistants have investigated a vast design space of scaffolding users’ interactions with LLMs during data analysis (Mcnutt et al., 2023). For example, prior works have explored providing phasewise or stepwise decomposition interfaces (Kazemitabaar et al., 2024a), structuring how users express intent and context with ephemeral UI (Cheng et al., 2024), manipulating intent (Kim, 2025), and embedding explanations and documentation support (Wang et al., 2022). These explorations brought different values to streamline the AI-assisted data analysis tasks, such as improving users’ perceived control (Kazemitabaar et al., 2024a) and lowering interaction barriers for novices (Wang et al., 2022; Cheng et al., 2024). However, they do not necessarily improve objective performance measures such as task success or completion time, as UI scaffolds on their own do not ensure learning outcomes or allow us to analyze the AI-use skills students need.
Beyond interface design, many efforts have explored mining logs of user interactions to derive useful insights. For example, recent works have explored logging and learning analytics in Jupyter notebooks to track students’ interactions with a chatbot AI (Valle Torre et al., 2025) or log novices’ data science workflow (Zhao et al., 2024). These approaches highlight the promise of leveraging naturally occurring clickstreams to study novice reasoning at scale (Claypool et al., 2001; Hijikata, 2004; Huang et al., 2012; Chen et al., 2025b), but also the challenges: raw logs are noisy, fragmented, and difficult to interpret without additional semantic structure e.g., a long pause may indicate either confusion or deep focus. To enable higher-level reasoning, we take inspiration from recent work, to extract more semantically meaningful insights from the raw logs by utilizing LLMs’ reasoning and Theory-of-Mind capabilities for reasoning process (Viswanathan et al., 2024; Rabinowitz et al., 2018; Kosinski, 2024; Strachan et al., 2024). To this end, we propose to combine log-based methods with behavior-grounded coding to identify success and failure patterns in AI-assisted notebook tasks, and identifying needed AI skills from the analysis. By semantically enriching logs into user story episodes and mapping them to concrete challenges, we aim to advance understanding of why some learners succeed while others fail, and what kinds of competencies need to be taught (Section 7.1, 7.2).
@@ 2.3. Training for AI-Assisted Data Analysis and Effective LLM Use
Prior works have also explored instructions specifically for AI-assisted data analysis tasks. For example, Zheng (2023) allowed ChatGPT use in a graduate data science task and found students reported that AI was more helpful for explaining code and less for open-ended problem-solving that needs context, especially on LLM-hard tasks where AI has limited capacities. LLM-enhanced feedback for an R-based data science tutoring system was also found to improve students’ performance, while the verbosity of AI explanations sometimes caused cognitive overload (Letteri and Vittorini, 2025). Nonetheless, these works mostly focus on analyzing surveys and prompt exchanges, without behavioral log data that is important for capturing the full experiences and interactions with LLM (Kazemitabaar et al., 2024b). Moreover, these studies may not represent the most authentic workflows of how students would use LLM for computational data analysis, when there are popular tools such as Google Colab (Google, 2025) where AI is embedded in the notebook environment.
@@
There have also been some works that focus on task-agnostic, domain-general AI-use skills, such as prompt engineering. However, students do not learn from code feedback and self-experimentation (Nguyen et al., 2024b); teaching specific prompt tips also may not be very effective at influencing students’ real prompting approaches and success, and also risks getting obsolete with model updates (Ma et al., 2025). Some prior works have shown promise in teaching requirement specification (Ma et al., 2025) or requirement elicitation (Lojo et al., 2025) for LLMs. However, no work has comprehensively mapped AI-use competencies across the full workflow or disentangled which skills are teachable via simple demonstrations versus guided practice and targeted feedback. Our study fills this gap by linking fine-grained behaviors to outcomes and by quantifying post-demo shifts in skill acquisition.
3. Course Structure and Task Setup
3.1. Course Context
Our IRB-approved study was conducted in Spring 2025 in Data Science for Product Managers (DSPM) at an R1 University in the United States. The course serves students with highly diverse academic and professional backgrounds, including product managers, engineers, and graduates of non-technical disciplines. This diversity creates a pedagogical tension: assignments that are trivial for some are overwhelming for others. The setting therefore provides an ideal opportunity to examine whether and how LLM support can reduce expertise gaps in authentic workflows.
Students worked in the Google Colab notebook with access to the built-in Gemini assistant (Google, 2025) [^1] We chose Google Colab instead of Jupyter Notebook because it is easier to set up for students who do not have any programming background, and Colab has an in-house LLM (Gemini) that is free to use and provides various affordances. Specific model used by Google Colab in Spring 2025 is unknown, but it could be gemini-2.0-flash or gemini-1.5-flash. We discuss the limitations of this approach compared to building on the open-source Jupyter Notebook in Section 8.2., but external GenAI systems such as ChatGPT were disallowed. Colab Gemini offered five interaction modes, including a chatbot, code cell suggestions, and buttons for error explanations, code explanations, and visualization generation with a dataframe (Figure 1 shows two of the main interfaces used for code generation: chatbot and code cell).
@@ 3.1. Course Context
Table 1. An overview and comparison of the homework setup, dataset, and rubrics. Overall task: analyze the data to gather some insights to present and guide product development across four stages of analysis (data cleaning, EDA, ML, storytelling).:
@@ 3.2. Homework Design and Procedure
Assignments were distributed as Google Colab notebook templates with minimal instructions and starter code, requiring students to “fill in the blank” using Python.
Students first completed a diagnostic HW0 during lecture to familiarize themselves with the tools and task sequence. HW0 covered all four tasks but used a different dataset than HW1-HW4 (details in Table 1). We asked students to spend about 15 minutes on each task, where half the class completed the first two or last two tasks without using GenAI, providing a baseline for comparison. Students then completed HW1–HW4 over the semester on a bi-weekly basis, each focused on a different aforementioned task. This setup enabled us to probe both short-term performance and longer-term workflow strategies, and to examine whether the LLM helps close the expertise gap. Details regarding the dataset, task, and rubrics are presented in Table 1.
For each assignment, students submitted (1) a completed notebook, (2) exported log files of interactions, and (3) a post-survey reflecting on challenges and AI usage (details in Section 4.2 and Appendix B). For HW0, students recorded their screen via Zoom, creating paired video–log traces for deep qualitative analysis. For HW1–HW4, bonus credit was offered for optional 60-minute think-alouds facilitated by one author, during which students would verbalize their thoughts while doing homework as usual.

### v2

# Not Everyone Wins with LLMs: Behavioral Patterns and Pedagogical Implications for AI Literacy in Programmatic Data Science

@@ Abstract.
LLMs promise to democratize technical work in complex domains like programmatic data analysis, but not everyone benefits equally. We study how students with varied experiences use LLMs to complete Python-based data analysis in computational notebooks in a graduate course. Drawing on homework logs, recordings, and surveys from 36 students, we ask: Which experience matters most, and how does it shape AI use? Our mixed-methods analysis shows that technical experience – not AI familiarity or communication skills – remains a significant predictor of success. Students also vary widely in how they leverage LLMs, struggling at stages of forming intent, expressing inputs, interpreting outputs, and assessing results. We identify success and failure behaviors, such as providing context or decomposing prompts, that distinguish effective use. These findings inform AI literacy interventions, highlighting that lightweight demonstrations improve surface fluency but are insufficient; deeper training and scaffolds are needed to cultivate resilient AI use skills.
large language model, AI-assisted data analysis, human-AI interaction, AI literacy
copyright: acmcopyrightjournalyear: 2026doi: 10.1145/3772318.3791283booktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26)price: 15.00isbn: 979-8-4007-2278-3/26/04ccs: Human-centered computing Interactive systems and toolsccs: Applied computing Computer-assisted instructionccs: Computing methodologies Natural language generation
@@ 1. Introduction
Large language models (LLMs) hold the promise of democratizing complex technical tasks such as data analysis by supporting programming, analysis, and modeling (Ma et al., 2025; Pickering et al., 2025). Despite the grand vision, evidence of who benefits and how remains mixed (Paradis et al., 2025). Prior work reports conflicting effects of AI agents on programming efficiency (Becker et al., 2025; Song et al., 2024); only some novices seem to perceive reduced barriers (Pickering et al., 2025; Kazemitabaar et al., 2023). In fact, performance often hinges on users’ prior experience, ranging from domain knowledge to AI literacy (Chen et al., 2024a; Nguyen et al., 2024; Zi et al., 2025). For instance, some work argues that less experienced developers benefit more from AI assistants (Liang et al., 2024; Peng et al., 2023), and others show that basic AI literacy helps users avoid common pitfalls in prompting or interpreting model outputs (Puppart and Aru, 2025; Annapureddy et al., 2025). However, these findings remain fragmented across different populations and tasks, making it challenging to weigh which types of experience matter most (e.g., whether we should focus on teaching domain knowledge or AI usage knowledge). Moreover, most prior analyses measure only outcomes (e.g., task success) and are conducted in controlled, short-term lab settings (Kazemitabaar et al., 2024a; Pickering et al., 2025; Zheng, 2023). As a result, we have yet to demystify how users’ AI interaction behaviors get shaped by different experiences, or the lack thereof, in real-world workflows.
@@
In this work, we examine how varied forms of experience influence AI use and outcomes in a naturalistic classroom setting (Section 3): a graduate-level course, Data Science for Product Management (DSPM). The class included a diverse cohort of 36 students from CS majors to business students transitioning into technical roles, allowing us to examine varying domain expertise, LLM literacy, and communication skills. Students independently completed data analysis assignments (e.g., Data Cleaning) in Google Colab with full access to the built-in Gemini LLM assistant, while their workflows were captured through logs, surveys, and screen recordings. Through this setup, we investigate three questions with a mix of quantitative and qualitative methods (Section 4):
RQ1 (Section 5): Do LLMs close the performance gap between students with different experience? We show that in open-ended, LLM-supported workflows, time constraints may remove the performance gap brought by prior technical experience. However, if given ample time, technical experience remains a significant predictor of success, and self-reported communication skills and LLM familiarity do not suffice. This suggests that students with stronger programming and data science backgrounds outperform their peers when using AI in real-world tasks.
@@
RQ2 (Section 6): How do experience differences translate to AI usage behaviors? Our analysis on over 7,000 log events shows that more technically experienced students use AI more strategically. They write clearer prompts and use AI proactively to improve plans (e.g., asking for better visualizations). In contrast, novices rely more on AI to resolve immediate challenges, such as debugging errors. These behavioral patterns offer a more nuanced explanation of why technical experience continues to matter in AI-supported environments.
@@
RQ3 (Section 7): What can we train students on AI usage? Comparing AI use behaviors before and after a lightweight instruction, we show that some AI-use skills (e.g., prompt quality) can be improved via live demonstration and extended time-on-task, while others (e.g., evaluation) likely require guided practice and feedback. We synthesize a set of AI-use competencies across Conceptual, Procedural, Metacognitive, and Dispositional knowledge dimensions.
Taken together, this work suggests that LLMs do not uniformly democratize programming but instead reshape when experience advantages appear. Our findings highlight AI literacy not as familiarity with tools, but as a set of transferable competencies for effective human–AI collaboration. We make three main contributions:
  • We provide empirical evidence that LLMs do not fully close the experience gap — while AI may lower immediate barriers under time pressure, sustained success still depends on technical skills.
  • We develop a behavior-oriented, LLM-based log annotation schema that segments logs into semantically meaningful user stories, and we open-source our code: https://github.com/mqo00/dspm
  • We identify concrete objectives and implications for future curricula and tool design to train effective AI collaborators.
@@ 2. Related Works
2.1. Novice vs. Expert AI Programming Strategy
HCI and educational research highlight systematic differences between novices and experts when collaborating with AI for programming. In agent-based modeling, experts could benefit more than novices and rely on stronger skills in task decomposition and evaluation of AI outputs (Chen et al., 2024a). In code comprehension, experts inspect the data more deeply, verify assumptions about transformations, and navigate outputs with clear goals — behaviors that novices underuse (Lum et al., 2025). Studies also document failure patterns such as “prompting rabbit holes,” where less-experienced users iterate on prompts without convergence (Tie et al., 2024). Beyond qualitative behavioral patterns, there has been mixed evidence regarding whether LLM use benefits people with various levels of expertise: some work found significant benefit from prior programming experience (Kazemitabaar et al., 2023), some find no significance (Pickering et al., 2025), while some found experts’ performance using AI is actually hindered (Becker et al., 2025). Additionally, prior studies focus on challenges brought by domain expertise differences, commonly measured by the level of prior experience in programming via self-reported surveys (Chen et al., 2024a, b; Lum et al., 2025), instead of analyzing different types of experiences. These findings motivate our work: measure whether LLMs reduce difference experience gaps, and analyze which behaviors influence AI use efficacy in authentic workflows.
@@ 2.2. Interfaces and Log Mining for AI-Assisted Data Analysis
AI-based computational notebook assistants have investigated a vast design space of scaffolding users’ interactions with LLMs during data analysis (Mcnutt et al., 2023; Gu et al., 2024). For example, prior works have explored providing phasewise or stepwise decomposition interfaces (Kazemitabaar et al., 2024a), structuring how users express intent and context with ephemeral UI (Cheng et al., 2024), manipulating intent (Kim, 2025), and embedding explanations and documentation support (Wang et al., 2022). These explorations brought different values to streamline the AI-assisted data analysis tasks, such as improving users’ perceived control (Kazemitabaar et al., 2024a) and lowering interaction barriers for novices (Wang et al., 2022; Cheng et al., 2024). However, they do not necessarily improve objective performance measures such as task success or completion time, as UI scaffolds on their own do not ensure learning outcomes or allow us to analyze the AI-use skills students need.
Beyond interface design, many efforts have explored mining logs of user interactions to derive useful insights. For example, recent works have explored logging and learning analytics in Jupyter notebooks to track students’ interactions with a chatbot AI (Valle Torre et al., 2025) or log novices’ data science workflow (Zhao et al., 2024). These approaches highlight the promise of leveraging naturally occurring clickstreams to study novice reasoning at scale (Claypool et al., 2001; Hijikata, 2004; Huang et al., 2012; Chen et al., 2025b), but also the challenges: raw logs are noisy, fragmented, and difficult to interpret without additional semantic structure, e.g., a long pause may indicate either confusion or deep focus. To enable higher-level reasoning, we extract more semantically meaningful insights from the raw logs by utilizing LLMs’ reasoning and Theory-of-Mind capabilities, inspired by recent work (Viswanathan et al., 2024; Rabinowitz et al., 2018; Kosinski, 2024; Strachan et al., 2024). We propose to combine log-based methods with behavior-grounded coding to identify success and failure patterns and identify needed AI skills. By semantically enriching logs into user story episodes, we aim to advance understanding of why some learners succeed while others fail, and what competencies need to be taught (Section 7).
@@
Screenshot of the Google Colab notebook environment showing the Gemini assistant. It displays code cells, a conversational sidebar, and buttons for explaining errors, explaining code, and generating visualizations.
@@ 2.3. Training for AI-Assisted Data Analysis and Effective LLM Use
Prior works have also explored instructions specifically for AI-assisted data analysis tasks. For example, Zheng (2023) allowed ChatGPT use in a graduate data science task and found that students reported that AI was more helpful for explaining code and less for open-ended problem-solving that needs context, especially on LLM-hard tasks where AI has limited capacities. LLM-enhanced feedback for an R-based data science tutoring system was also found to improve students’ performance, while the verbosity of AI explanations sometimes caused cognitive overload (Letteri and Vittorini, 2025). Nonetheless, these works mostly focus on analyzing surveys and prompt exchanges, leaving out behavioral log data that is important for capturing the full experiences and interactions with LLM (Kazemitabaar et al., 2024b). Moreover, these studies may not represent the most authentic workflows of how students would use LLM for computational data analysis, when there are popular tools such as Google Colab (Google, 2025) where AI is embedded in the notebook environment.
Similar gaps exist in the broader LLM for CS education literature, where classroom-based studies, especially in graduate-level as in our context, remain limited (Stone, 2024; Raihan et al., 2025). In parallel, prior work primarily extracts student workflow or prompting patterns with surveys (Gorson Benario et al., 2025; Scholl and Kiesler, 2024), interviews (Tie et al., 2024), or think-aloud (Lum et al., 2025), overlooking analysis on log data. Log data provides fine-grained traces of user behavior that reveal naturalistic actions without interference and helps us analyze interaction strategies linked to performance, which are not accessible through self-report, grades, or observation (Dumais et al., 2014; Ihantola et al., 2015).
@@
There have also been some works that focus on task-agnostic, domain-general AI-use skills, such as prompt engineering. However, students do not learn from code feedback and self-experimentation (Nguyen et al., 2024). Teaching specific prompt tips may not be very effective at influencing students’ real success beyond perceived quality (Sawalha et al., 2024), and risks getting obsolete with model updates (Ma et al., 2025). Some prior works have shown promise in teaching requirement specification (Ma et al., 2025) or requirement elicitation (Lojo et al., 2025) for LLMs. However, no work has comprehensively mapped AI-use competencies across the full workflow or disentangled which skills are harder to teach. Our study fills this gap by linking fine-grained behaviors to outcomes and quantifying post-demo shifts in skill acquisition.
@@
Table 1. An overview and comparison of the homework setup, dataset, and rubrics. Overall task: analyze the data to gather some insights to present and guide product development across four stages of analysis (data cleaning, EDA, ML, storytelling).
@@
Overview of homework setup, datasets, and grading rubrics. HW0 served as a diagnostic with short, timed tasks, while HW1–HW4 were bi-weekly assignments with extended time. Rubrics for each task (data cleaning, EDA, ML, storytelling) evaluated correctness and completeness on a 0–3 scale.
3. Course Structure and Task Setup
3.1. Course Context
Our IRB-approved study was conducted in Spring 2025 in Data Science for Product Managers (DSPM) at an R1 University in the United States. The course serves students with highly diverse academic and professional backgrounds, including product managers, engineers, and graduates of non-technical disciplines. This diversity creates a pedagogical tension: assignments that are trivial for some are overwhelming for others. The setting, therefore, provides an ideal opportunity to examine whether and how LLM support can reduce experience gaps in authentic workflows.
Students worked in the Google Colab notebook with access to the built-in Gemini assistant (Google, 2025) [^1] We chose Google Colab instead of Jupyter Notebook because it is easier to set up for students who do not have any programming background, and Colab has an in-house LLM (Gemini) that is free to use and provides various affordances. Specific model used by Google Colab in Spring 2025 is unknown, but it could be gemini-2.0-flash or gemini-1.5-flash. We discuss the limitations of this approach compared to building on the open-source Jupyter Notebook in Section 8.2., but external GenAI systems such as ChatGPT were disallowed. Colab Gemini offered five interaction modes, including a chatbot, code cell suggestions, and buttons for error explanations, code explanations, and visualization generation with a dataframe (Fig. 1 shows two of the main interfaces used for code generation: chatbot and code cell).
@@ 3.2. Homework Design and Procedure
Assignments were distributed as Google Colab notebook templates with minimal instructions and starter code, requiring students to “fill in the blank” using Python. We used unpopular datasets that are further mutated or collected via crowdsourcing in the prior semester, so solutions cannot be found via internet searches. We verified that the assignment could not be trivially completed by Colab Gemini or ChatGPT, by importing the template and dataset and prompting the models to complete all tasks with intuitive instructions.
Students first completed a diagnostic HW0 during lecture to familiarize themselves with the tools and task sequence. HW0 covered all four tasks but used a different dataset than HW1-HW4 (described in Table 1). We asked students to spend about 15 minutes on each task, where half the class completed the first two or last two tasks without using GenAI, providing a baseline for comparison. Students then completed HW1–HW4 over the semester on a bi-weekly basis, each focused on a different aforementioned task. This setup enabled us to probe both short-term performance and longer-term workflow strategies, and to examine whether the LLM helps close the experience gap. Details regarding the dataset, task, and rubrics are presented in Table 1.
For each assignment, students submitted (1) a completed notebook, (2) exported log files of interactions, and (3) a post-survey reflecting on challenges and AI usage (details in Section 4.2 and Appendix C). For HW0, students recorded their screen via Zoom, creating paired video–log traces for deep qualitative analysis. For HW1–HW4, bonus credit was offered for optional 60-minute think-alouds facilitated by one author, during which students would verbalize their thoughts while doing homework as usual.
@@
3.3. Instructional Baseline Demo
Between HW0 and HW1, the instructor provided a 90-minute live-coding demonstration of the HW0 data cleaning task. The demonstration showcased examples of AI use behaviors in the codebook (Table 2), with explanations of when and why each behavior could be useful. For example, the instructor explained “If you find unfamiliar concepts, you can actually ask Gemini to get you started. We can ask, What are some key steps in Data cleaning?” while typing and sending the actual question and walking through Gemini’s answer. The demo also covered the task environment (e.g., overview of Colab as a product) as well as the AI context space (e.g., the code cell generation has access to the entire notebook, while the chatbot sidebar does not have access to code cell outputs). Students could re-watch the recording, too. The demo’s lecturing nature has limitations, as it can be less cognitively engaging than active learning and may create an illusion of fluency without fostering durable understanding (Chi and Wylie, 2014; Deslauriers et al., 2019). However, this is a lightweight instruction baseline that enables analysis for what AI use competency needs to be taught with greater depth (RQ3).